### v3

# FitCF: A Framework for Automatic Feature Importance-guided Counterfactual Example Generation

Abstract
Counterfactual examples are widely used in natural language processing (NLP) as valuable data to improve models, and in explainable artificial intelligence (XAI) to understand model behavior. The automated generation of counterfactual examples remains a challenging task even for large language models (LLMs), despite their impressive performance on many tasks. In this paper, we first introduce ZeroCF, a faithful approach for leveraging important words derived from feature attribution methods to generate counterfactual examples in a zero-shot setting. Second, we present a new framework, FitCF [^1] Code is available at: https://github.com/qiaw99/FitCF. FitCF stands for “Feature Importance-guided Counterfactual Example Generation”., which further verifies aforementioned counterfactuals by label flip verification and then inserts them as demonstrations for few-shot prompting, outperforming three state-of-the-art baselines. Through ablation studies, we identify the importance of each of FitCF’s core components in improving the quality of counterfactuals, as assessed through flip rate, perplexity, and similarity measures. Furthermore, we show the effectiveness of LIME and Integrated Gradients as backbone attribution methods for FitCF and find that the number of demonstrations has the largest effect on performance. Finally, we reveal a strong correlation between the faithfulness of feature attribution scores and the quality of generated counterfactuals, which we hope will serve as an important finding for future research in this direction.
FitCF: A Framework for Automatic Feature Importance-guided Counterfactual Example Generation
Qianli Wang1,2,22footnotemark: 2   Nils Feldhus1,2,6   Simon Ostermann2,4,5 Luis Felipe Villa-Arenas1,2,3   Sebastian Möller1,2   Vera Schmitt1,2 1Quality and Usability Lab, Technische Universität Berlin 2German Research Center for Artificial Intelligence (DFKI)  3Deutsche Telekom 4Saarland Informatics Campus  5Centre for European Research in Trusted AI (CERTAIN) 6BIFOLD – Berlin Institute for the Foundations of Learning and Data 22footnotemark: 2 Correspondence: qianli.wang@tu-berlin.de
Figure 1: Given an instance $x$ from the AG News dataset classified as “sci/tech”, our ZeroCF approach generates few-shot examples, whose important words are determined by LIME for a BERT model. FitCF then generates a counterfactual $\tilde{x}$ on this basis. The edits to original instance $x$ are highlighted in orange, yielding $\tilde{x}$ which is classified as “business”.
1 Introduction
The advent of increasingly complex and opaque LLMs has triggered a critical need for explainability and interpretability of such models. Counterfactuals, which are minimally edited inputs that yield different predictions compared to reference inputs Miller (2019); Ross et al. (2021); Madsen et al. (2022) are widely used in XAI and NLP. Applications include creating new data points for improving models in terms of performance Kaushik et al. (2020); Sachdeva et al. (2024) and robustness Gardner et al. (2020); Ross et al. (2021), and understanding the black-box nature of models Wu et al. (2021); Wang et al. (2024a, b). Crowd-sourcing counterfactuals can be costly, inefficient, and impractical Chen et al. (2023), particularly in a specialized domain such as medicine. LLM-based counterfactual generation offers a more efficient and scalable alternative. Despite advancements in counterfactual generation techniques and the demonstrated versatility of LLMs across tasks Wu et al. (2021); Bhan et al. (2023a); Li et al. (2024), the efficacy of LLMs in producing high-quality counterfactuals in a zero-shot setting, as well as the effective construction of valid counterfactuals as demonstrations to enable few-shot prompting, remains an open question Bhattacharjee et al. (2024a). Additionally, the combination of widely used interpretability methods, with the goal to exploit their combined benefits, has been insufficiently explored within XAI research Treviso et al. (2023); Baeumel et al. (2023); Bhan et al. (2023a).
To this end, we first present ZeroCF, a method to combine feature importance with counterfactual generation by leveraging important words identified through feature attribution scores for a fine-tuned BERT Devlin et al. (2019) on the target dataset, evaluated on four representative feature importance methods (§4.4). The generation of counterfactuals with ZeroCF is performed by prompting LLMs with extracted important words in a zero-shot setting without any auxiliary counterfactual data (§3.1). We then propose the FitCF framework (Figure 1), which uses ZeroCF-generated counterfactuals following a label flip verification step as demonstrations for few-shot prompting without relying on human-crafted examples (§3.2).
Secondly, we evaluate ZeroCF and FitCF on two NLP tasks - news topic classification and sentiment analysis - using three baselines, Polyjuice Wu et al. (2021), BAE Garg and Ramakrishnan (2020) and FIZLE Bhattacharjee et al. (2024a). The automatic evaluation employs three automated metrics: Label flip rate, fluency, and edit distance. Both ZeroCF and FitCF significantly outperform Polyjuice and BAE, with ZeroCF surpassing FIZLE in most cases and FitCF consistently exceeding three state-of-the-art baselines and ZeroCF.
Thirdly, we perform ablation studies on three key components of FitCF: (1) Important words; (2) the number of demonstrations; (3) label flip verification. The results reveal that all three components contribute positively to improving the quality of counterfactuals, as measured by label flip rate, fluency, and edit distance, with the number of demonstrations being the most influential. In addition, FitCF exhibits greater robustness and achieves higher overall quality when combined with LIME and SHAP compared to its combination with Gradient and Integrated Gradients.
Lastly, we conduct a correlation analysis between the quality of generated counterfactuals and the faithfulness of feature attribution scores as used in ZeroCF and FitCF. The analysis reveals that LIME and SHAP can produce more faithful feature attribution scores compared to Gradient and Integrated Gradients. Furthermore, we observe a strong correlation between the faithfulness of these feature attribution scores and the quality of counterfactuals generated by FitCF.
Figure 2: The upper part of the figure illustrates how counterfactuals are generated by ZeroCF using important words extracted by the explainer (BERT) through various feature important methods (Gradient, Integrated Gradients, LIME, SHAP). Lower part of the figure shows the pipeline of FitCF involving demonstrations selection, automatic construction of counterfactual examples by ZeroCF, label flip verification, and counterfactual generation.
2 Related Work
Counterfactual Generation
MICE generates contrastive edits that change the prediction to a given contrast prediction Ross et al. (2021). Polyjuice uses a fine-tuned GPT-2 Radford et al. (2019) to specify the type of edit needed to generate counterfactual examples Wu et al. (2021). DISCO Chen et al. (2023) uses the GPT-3 fill-in-the-blank mode, which is not available in most open-source LLMs Chen et al. (2023). Tigtec Bhan et al. (2023b) utilizes local feature importance to identify words that significantly influence a model’s prediction and masks these important words using a fine-tuned masking model. Bhattacharjee et al. (2024b) identify the latent features in the input text and the input features associated with the latent features to generate counterfactual examples, which is criticized due to the additional level of complexity with no significant performance gain Delaunay et al. (2024). FIZLE Bhattacharjee et al. (2024a) shares the most similarity with FitCF and uses LLMs as pseudo-oracles to generate counterfactuals with the assistance of LLM-generated important words in a zero-shot setting.
Combination of Interpretability Methods
Recent works have explored the possibility to combine different XAI methods. Wang et al. (2021) propose a feature importance-aware attack, which disrupts important features that consistently influence the model’s decisions. Gressel et al. (2023) identify perturbations in the feature space to produce evasion attacks. Treviso et al. (2023) present the framework, CREST, to generate counterfactual examples by combining rationalization with span-level masked language modeling. Krishna et al. (2023) employ various post-hoc explanations for rationalization, extending beyond counterfactuals, in contrast to CREST. Bhan et al. (2023a) propose a method to determine impactful input tokens with respect to generated counterfactual examples. In contrast, FitCF uses feature importance to guide counterfactual example generation.
3 Methodology
3.1 ZeroCF
Bhattacharjee et al. (2024a) introduced FIZLE, which generates counterfactuals in a zero-shot setting by prompting the LLM with important words identified by the LLM itself. However, these extracted words may be unfaithful or hallucinated Li et al. (2023) [^2] Applying Llama3-8B with FIZLE on AG News, we find that for 64.5% of the instances, a subset of generated important words is hallucinated, i.e., absent from the original input.. To address this limitation, we propose ZeroCF (Figure 2; examples are provided in Table 7), which relies on the most attributed words based on feature attribution scores determined by various explanation methods for the predictions of a BERT model fine-tuned on the target dataset. Feature importance involves determining how significant an input feature is for a given output Madsen et al. (2022), which we find to enhance the counterfactual generation process (§6.1).
Prediction
Given an input $x$ from the dataset $\mathcal{D}$, we leverage a BERT model $\mathcal{M_{D}}$ fine-tuned on $\mathcal{D}$ [^3] Detailed information, e.g., accuracy, about the deployed BERT models is provided in Appendix B. to obtain the prediction $y_{pred}$ for the given input $x$:
$y_{pred}=\mathcal{M_{D}}(x)$ (1)
Feature Attribution Scores
Then we deploy an explainer $\mathcal{E}$ with access to the model $\mathcal{M_{D}}$, which employs various feature importance methods $f$4.4) to acquire feature attributions scores $s$ based on the prediction $y_{pred}$ and the given input $x$:
$s=\mathcal{E}(x,y_{pred},f,\mathcal{M_{D}})$ (2)
Counterfactual Generation
Finally, we identify the top-attributed words [^4] The top attributed words are further post-processed by replacing the “[CLS]” and “[SEP]” special tokens if any, with the subsequent attributed words and by merging tokenized subwords if one of them is a top attributed word. $w$ based on feature attribution scores $s$ and deploy an LLM $\mathcal{L}$ in a zero-shot setting to generate the counterfactual $\tilde{x}$ with the prompt $p$A.1), which consists of task instruction $i$, words $w$, the prediction $y_{pred}$, and the input $x$:
$\tilde{x}=\mathcal{L}(p)$ (3)
3.2 FitCF
While ZeroCF mitigates the issue of hallucinated important words extracted by the LLM, the counterfactuals generated by ZeroCF may fail to flip the prediction, e.g., due to the limited capability of zero-shot prompting Brown et al. (2020). To address it, we propose FitCF (Figure 1, Figure 2), inspired by Auto-CoT Zhang et al. (2023), which combines two interpretability methods, feature importance and counterfactual examples, leveraging their complementary advantages and automatically constructs demonstrations by ZeroCF incorporating label-flip verification. Verified demonstrations subsequently enable few-shot prompting in FitCF.
top-$k$ Examples Sampling
In order to diversify demonstration selection An et al. (2023); Zhang et al. (2023) and construct demonstrations automatically, we first convert all instances from the dataset $\mathcal{D}$ into sentence embeddings using SBERT [^5] https://huggingface.co/sentence-transformers/all-mpnet-base-v2, and then apply $k$-means clustering on these sentence embeddings to form $k$ clusters [^6] Clustering visualizations are given in Appendix C., where $k$ does not necessarily correspond to the exact number of predefined dataset labels. Afterwards, we select a total of $c$ instances which are closest to the centroid of each cluster [^7] Selected examples and their corresponding counterfactuals for a given instance are provided in Appendix D.. In such a way, we diversify the demonstrations, potentially mitigating any misleading effects caused by ZeroCF, which may produce flawed counterfactuals. Finally, ZeroCF is employed to generate counterfactuals for the $c$ selected instances using simple heuristics.
Label Flip Verification
Subsequently, in order to validate the generated counterfactuals and to prevent incorrect counterfactuals from misleading the LLM Turpin et al. (2023), we employ the same BERT model $\mathcal{M_{D}}$3.1) to make predictions on $c$ generated counterfactuals $\mathcal{C}=\{\tilde{x}_{1},\tilde{x}_{2},...,\tilde{x}_{c}\}$ and the original input $\mathcal{X}=\{x_{1},x_{2},...,x_{c}\}$ individually and assess whether the labels are inconsistent:
$\forall i\in\{1,2,\cdots,c\}:\hat{y}_{x_{i}}=\mathcal{M_{D}}(x_{i})$ (4)
$\forall i\in\{1,2,\cdots,c\}:\hat{y}_{\tilde{x}_{i}}=\mathcal{M_{D}}(\tilde{x}_{i})$ (5)
The generated counterfactuals $\tilde{x}_{i}$, where the predicted labels remain consistent $\hat{y}_{\tilde{x}_{i}}=\hat{y}_{x_{i}}$, are excluded from the demonstrations for further process to ensure the validity of the generated counterfactuals. In the end, we obtain $m$ counterfactuals, where $m\leq c$. To maintain a consistent number of demonstrations ($\ell$) for each input, if $m<\ell$, additional examples are iteratively selected based on their proximity to the cluster centroid, until the required number of demonstrations is achieved. Moreover, given we select $\ell$ demonstrations with $\ell>>k$, the influence of the number of clusters, $k$, on the final performance is diminished.
Counterfactual Generation
For a given input $x$, $\ell$ input-counterfactual pairs generated by ZeroCF are used as demonstrations, along with important words $w$ extracted based on the feature attribution scores $s$ generated by BERT3.1), to prompt the LLM to generate the counterfactual for the input $x$ in a few-shot setting (Figure 2, §A.2).
3.3 Considerations for Choice of Models
In ZeroCF, feature attributions are generated for a BERT model’s predictions, based on which important words are then extracted(§3.1). Moreover, in FitCF, the same BERT model serves as a label flip verifier (§3.2). BERT is chosen as our design choice, because it is a representative encoder language model with a strong efficiency-performance balance. We emphasize that any model capable of performing classification tasks effectively can be used as a label flip verifier or for generating feature attribution scores. For encoder-only architectures like the BERT model employed in our study, tools like Ferret Attanasio et al. (2023) can be used to derive feature attribution scores (§4.4). For encoder-decoder or decoder-only architectures, tools like Inseq Sarti et al. (2023) can generate such scores.
4 Experimental Setup
4.1 Baselines
We employ three approaches as baselines for ZeroCF and FitCF, including FIZLE that achieves state-of-the-art performance in zero-shot counterfactual generation, which we aim to enhance.
BAE
BAE is an adversarial attack method that employs BERT to perturb input text by replacing masked words Garg and Ramakrishnan (2020).
Polyjuice
Polyjuice allows users to control perturbation types and deploys a GPT-2888Although Polyjuice utilizes a relatively small model, GPT-2, for generating counterfactuals, and BAE uses BERT to replace words based on embeddings, we consider both of them suitable baseline methods for FitCF. This is because the deployed GPT-2 is fine-tuned on a counterfactual example dataset. Furthermore, FIZLE relies on zero-shot prompting and achieves state-of-the-art performance. to generate counterfactuals by framing the task as a conditional text generation problem Wu et al. (2021).
FIZLE
FIZLE employs an LLM to identify important words and prompts the LLM with these words in a zero-shot setting to generate counterfactuals Bhattacharjee et al. (2024a).
4.2 Dataset
Following Nguyen et al. (2024); Bhattacharjee et al. (2024a), we demonstrate the validity of ZeroCF and FitCF by applying them to two NLP tasks: News topic classification and sentiment analysis [^9] Details on label distributions and example instances from the datasets used can be found in Appendix E..
AG News
AG News Zhang et al. (2015) contains news articles created by combining the titles and description fields of articles from four categories: World, Sports, Business, and Sci/Tech.
SST2
SST2 Socher et al. (2013) is part of the larger Stanford Sentiment Treebank and focuses specifically on binary sentiment classification of natural language movie reviews. Each sentence is labeled as either negative or positive.
4.3 Models for Counterfactual generation
We select three open source state-of-the-art instruction fine-tuned LLMs with increasing parameter sizes [^10] More details about deployed models and inference time are provided in Appendix F.: Llama3-8B AI@Meta (2024), and Qwen2.5-{32B,72B} Team (2024).
4.4 Feature Importance
Ferret Attanasio et al. (2023) is a framework that provides post-hoc explanations for LLMs and can evaluate these explanations based on faithfulness and plausibility. We use ferret to generate feature attribution scores, selecting the following feature importance methods $f$: Gradient Simonyan et al. (2014), LIME Ribeiro et al. (2016), Integrated Gradients Sundararajan et al. (2017), and SHAP Lundberg and Lee (2017).
5 Evaluation
5.1 Automatic Evaluation
The generated counterfactuals are evaluated using the following three automated metrics.
Soft Label Flip Rate
The Soft Label Flip Rate (SLFR) measures the frequency at which newly perturbed examples alter the original label to a different label Ge et al. (2021); Nguyen et al. (2024); Bhattacharjee et al. (2024b). For a dataset with $N$ instances, we calculate SLFR as follows:
$SLFR=\frac{1}{N}\sum_{n=1}^{N}\mathds{1}(y_{k}^{{}^{\prime}}\neq y_{k})$
where $\mathds{1}$ is the indicator function, $y_{k}$ is the original label and $y_{k}^{{}^{\prime}}$ is the predicted label after the perturbation. Note that we use the same LLM for both counterfactual generation and classification [^11] The accuracy and error rate of the deployed LLMs, along with the prompt instruction used are provided in Appendix G..
Perplexity
Perplexity (PPL) is defined as the exponential of the average negative log-likelihood of a sequence. PPL can measure the naturalness of the text distribution and how fluently the model can output the next word given the previous words Fan et al. (2018). Given a sequence $X=(x_{0},x_{1},\cdots,x_{t})$, PPL of $X$ is calculated as:
$PPL(X)=\exp\left\{\frac{1}{t}\sum_{i}^{t}\log{p_{\theta}(x_{i}|x_{<i})}\right\}$
Following Wang et al. (2023); Nguyen et al. (2024); Bhattacharjee et al. (2024a), we deploy GPT-2 to calculate PPL in our experiments due to its proven effectiveness in capturing such text distributions.
Textual Similarity (TS)
The counterfactual $\tilde{x}$ should be as similar as the original input $x$ Madaan et al. (2021), where lower distances indicate greater similarity. We use normalized word-level Levenshtein distances $d$ to capture all edits, which is widely used by the research community Ross et al. (2021); Treviso et al. (2023):
$TS=\frac{1}{N}\sum_{i=1}^{N}\frac{d(x_{i},\tilde{x}_{i})}{|x_{i}|}$
5.2 Ablation Study
As illustrated in Figure 2, FitCF comprises three core components: Important words; demonstrations; and label flip verification. Accordingly, we conduct a comprehensive ablation study to evaluate the importance of each component individually. The experiments are conducted using Qwen2.5-72B, as Qwen2.5-72B particularly struggles to generate high-quality counterfactual examples compared to Llama3-8B and Qwen2.5-32B (Table 1, Table 3).
5.2.1 Effect of Important Words
To assess the contribution of important words identified by BERT using different feature importance methods to counterfactual generation, we conduct the experiment using FitCF omitting any pre-identified important words.
Approach Dataset AG News ($PPL=95.72$) SST2 ($PPL=309.53$)
Model Method SLFR $\uparrow$ PPL $\downarrow$ TS $\downarrow$ SLFR $\uparrow$ PPL $\downarrow$ TS $\downarrow$
Polyjuice GPT2 - 18.60% 121.76 0.50 29.00% 258.32 0.71
BAE BERT - 19.50% 168.44 0.12 47.00% 367.06 0.09
FIZLE Llama3-8B - 93.50% 123.67 0.61 95.50% 202.22 0.52
Qwen2.5-32B - 49.00% 53.07 1.14 86.80% 167.51 0.66
Qwen2.5-72B - 21.50% 84.09 0.22 92.00% 257.91 0.43
ZeroCF (Ours) Llama3-8B Gradient 93.50% 102.56 0.38 97.50% 239.15 0.46
IG 95.50% 109.09 0.27 99.50% 222.51 0.42
LIME 97.50% 107.72 0.39 97.00% 264.91 0.42
SHAP 98.00% 99.08 0.27 94.00% 204.76 0.46
Qwen2.5-32B Gradient 68.00% 62.63 2.10 70.50% 205.06 0.48
IG 51.00% 60.45 0.76 91.00% 222.57 0.64
LIME 56.00% 63.75 0.84 90.50% 576.59 0.62
SHAP 55.50% 61.68 0.79 93.00% 191.00 0.60
Qwen2.5-72B Gradient 16.67% 74.19 0.21 88.50% 263.47 0.34
IG 24.50% 92.47 0.22 92.00% 281.10 0.46
LIME 23.00% 72.73 0.71 85.00% 289.20 0.30
SHAP 25.00% 73.92 0.74 86.50% 319.60 0.22
FitCF (Ours) Llama3-8B Gradient 94.50% 86.90 0.21 99.80% 159.57 0.47
IG 96.00% 87.67 0.23 100.00% 161.88 0.48
LIME 95.50% 75.15 0.19 100.00% 151.22 0.48
SHAP 94.00% 260.57 0.21 100.00% 157.36 0.49
Qwen2.5-32B Gradient 56.00% 62.97 0.73 89.00% 214.25 0.51
IG 57.50% 57.01 0.68 90.50% 221.64 0.49
LIME 56.00% 57.45 0.79 89.50% 174.34 0.52
SHAP 62.00% 57.64 0.78 89.50% 157.09 0.52
Qwen2.5-72B Gradient 77.00% 62.13 0.99 96.00% 595.71 0.38
IG 42.00% 63.54 0.33 95.00% 207.55 0.39
LIME 45.00% 61.54 0.35 96.50% 240.94 0.41
SHAP 38.96% 67.28 0.34 96.50% 590.94 0.39
Table 1: Automatic evaluation results of counterfactuals generated by Polyjuice, BAE, FIZLE, ZeroCF, and FitCF with Llama3-8B, Qwen2.5-32B, and Qwen2.5-72B using Soft Label Flip Rate (SLFR), Perplexity (PPL), and Textual Similarity (TS) on AG News and SST2. Bold faced values indicate for each approach, which feature importance method is the best performing according to the respective metric.
5.2.2 Effect of Number of Demonstrations
In FitCF, as $k$ clusters are obtained through clustering, and due to the difficulty and complexity of counterfactual example generation, we set the number of demonstrations to twice the number of clusters for each dataset ($\ell=2k$; §3.2), which results in 10 demonstrations for AG News and 8 for SST2, respectively (Figure 4). To examine the effect of the number of demonstrations and assess the necessity of doubling the number of demonstrations to $2k$, we further evaluate the quality of counterfactual examples generated by FitCF, with the number of demonstrations set to the number of clusters ($\ell=k$).
5.2.3 Effect of Label Flip Verification
To ensure the validity of the selected demonstrations and prevent incorrect examples from misleading the LLM Rubin et al. (2022); Turpin et al. (2023), FitCF incorporates a label flip verifier (§3.2). This verifier is implemented using a fine-tuned BERT model (Table 6) trained on the target dataset. To assess the impact of label flip verification, we conduct an ablation study by excluding label flip verification for comparative analysis.
5.3 Correlation Analysis
As we deploy various feature importance methods to generate counterfactuals synergistically (Figure 2), which can then be applied as demonstrations in FitCF, we investigate the correlation between the quality of the feature attribution scores and the quality of generated counterfactuals. The feature attribution scores are evaluated based on faithfulness using ferret Attanasio et al. (2023). For faithfulness evaluation, we employ three metrics: comprehensiveness, sufficiency DeYoung et al. (2020) and Kendall’s $\tau$ correlation with Leave-One-Out token removal Jain and Wallace (2019).
Dataset Method SLFR PPL TS
AG News Gradient 41.50% ($\downarrow$35.50%) 67.85 ($\downarrow$5.72) 0.36 ($\uparrow$0.63)
IG 37.50% ($\downarrow$4.50%) 67.85 ($\downarrow$4.31) 0.37 ($\uparrow$0.62)
LIME 40.68% ($\downarrow$4.32%) 66.08 ($\downarrow$2.54) 0.35 ($\uparrow$0.02)
SHAP 37.00% ($\downarrow$1.96%) 84.14 ($\downarrow$16.86) 0.51 ($\downarrow$ 0.17)
SST2 Gradient 93.50% ($\downarrow$2.50%) 214.27 ($\uparrow$381.44) 0.42 ($\downarrow$ 0.04)
IG 95.00% (- 0.00%) 214.27 ($\downarrow$6.72) 0.42 ($\downarrow$0.02)
LIME 95.50% ($\downarrow$1.00%) 278.78 ($\downarrow$37.84) 0.41 (-0.00)
SHAP 96.00% ($\downarrow$0.50%) 290.57 ($\uparrow$-300.37) 0.43 ($\downarrow$0.04)
Table 2: Automatic evaluation results of counterfactuals generated by FitCF using Qwen2.5-72B, with demonstrations generated by ZeroCF without specifying important words.
6 Results
6.1 Automatic Evaluation
Table 1 demonstrates that our proposed approaches, ZeroCF and FitCF, consistently outperform Polyjuice and BAE easily, which exhibit relatively low SLFR. Notably, BAE achieves the lowest edit distance compared to Polyjuice, FIZLE, ZeroCF, and FitCF, as it only replaces masked words based on textual embeddings. For AG News dataset using Qwen2.5-32B, the edit distance is comparatively higher than that of Polyjuice, and the other baseline, FIZLE, also shows a larger edit distance compared to Polyjuice. For SST2 dataset, Qwen2.5-72B tends to generate counterfactuals that are less natural and fluent when leveraging ZeroCF and FitCF. Interestingly, Llama3-8B, the smallest model among all evaluated LLMs, achieves the best overall performance. In contrast, Qwen2.5-72B generally underperforms compared to both Llama3-8B and Qwen2.5-32B, as Qwen2.5-72B has a stronger capability to discern the underlying context, making it less prone to flipping labels (Appendix D).
Additionally, we observe that ZeroCF does not outperform FIZLE in some cases, e.g., with Qwen2.5-72B on SST2 dataset. However, in most cases, ZeroCF offers noticeable advantages in enhancing the quality of counterfactuals compared to FIZLE. Furthermore, we find that Integrated Gradients and SHAP contribute more positively to the quality of counterfactuals, on average [^12] We do not consider the number of times a feature importance method achieves the maximum value in tables, but rather the average ranking of a method across all datasets., compared to other feature importance methods.
Importantly, FitCF emerges as the most effective method for generating high-quality counterfactuals, consistently outperforming three baselines and ZeroCF across all evaluated settings, underscoring its robustness and effectiveness. This demonstrates the advantage of combining feature importance with the counterfactual generation process. Under the FitCF framework, Integrated Gradients and LIME illustrate superior performance in generating counterfactuals compared to the other two approaches.
6.2 Ablation Study
The results of the ablation studies are presented in Table 2, 3, 4, where for PPL and TS, an upward (downward) arrow signifies that a decrease (increase) in the value corresponds to an improvement (deterioration) in both metrics.
6.2.1 Effect of Important Words
Table 2 shows that for AG News, SLFR decreases across all methods, with the most significant decline observed when using Gradient. Concurrently, PPL improves and edit distances generally increases, suggesting that the generated counterfactuals diverge more from the original text, except when using SHAP. In contrast, for SST2, SLFR remains consistently high, with slight decreases. PPL exhibited mixed results, with both notable increases and decreases depending on the method, reflecting variability in fluency. Meanwhile, edit distance either decreases or remains unchanged. Overall, FitCF with SHAP demonstrates the highest robustness when important words are not specified, whereas Gradient is particularly sensitive to the inclusion of important words.
Dataset Method SLFR PPL TS
AG News Gradient 13.50% ($\downarrow$63.50%) 66.74 ($\downarrow$4.61) 0.27 ($\uparrow$0.72)
IG 15.50% ($\downarrow$22.00 %) 64.28 ($\downarrow$0.74) 0.27 ($\uparrow$0.06)
LIME 18.00% ($\downarrow$27.00%) 68.28 ($\downarrow$6.74) 0.27 ($\downarrow$0.08)
SHAP 14.00% ($\downarrow$24.96%) 64.06 ($\uparrow$3.22) 0.28 ($\uparrow$0.06)
SST2 Gradient 89.00% ($\downarrow$7.00%) 235.08 ($\uparrow$360.63) 0.36 ($\uparrow$0.02)
IG 93.50% ($\downarrow$1.50 %) 266.09 ($\downarrow$58.54) 0.39 (-0.00)
LIME 91.50% ($\downarrow$5.00%) 250.70 ($\downarrow$9.76) 0.39 ($\uparrow$0.02)
SHAP 92.00% ($\downarrow$4.50%) 583.42 ($\uparrow$7.52) 0.38 ($\uparrow$0.01)
Table 3: Automatic evaluation results of counterfactuals generated by FitCF with Qwen2.5-72B using $k$ demonstrations.
6.2.2 Effect of Number of Demonstrations
As shown in Table 3, we find that the number of demonstrations plays an critical role in the performance of FitCF. For AG News, SLFR declines precipitously when the number of clusters ($k$) is used as the number of demonstrations (§5.2.2), while the edit distance shows a slight improvement. In comparison, for SST2, the degree of SLFR diminishment is less conspicuous.
Furthermore, Table 3 reveals that in general, FitCF with Integrated Gradients and SHAP exhibits greater robustness compared to Gradient and LIME. In particular, FitCF with Gradient demonstrates the highest sensitivity, with a strong decline in quality as the number of demonstrations decreases.
6.2.3 Effect of Label Flip Verification
Dataset Method SLFR PPL TS
AG News Gradient 34.00% ($\downarrow$43.00%) 63.27 ($\downarrow$1.14) 0.33 ($\uparrow$ 0.66)
IG 40.50% ($\downarrow$1.50%) 64.65 ($\downarrow$1.11) 0.35 ($\downarrow$ 0.02)
LIME 42.50% ($\downarrow$2.50%) 65.23 ($\downarrow$3.69) 0.35 (- 0.00)
SHAP 34.00% ($\downarrow$4.96%) 65.30 ($\uparrow$1.98) 0.34 (- 0.00)
SST2 Gradient 94.50% ($\downarrow$1.50%) 222.52 ($\uparrow$373.19) 0.36 ($\uparrow$0.02)
IG 94.50% ($\downarrow$2.00%) 240.11 ($\downarrow$32.56) 0.39 (- 0.00)
LIME 96.00% ($\downarrow$0.50%) 245.79 ($\downarrow$4.85) 0.40 ($\uparrow$0.01)
SHAP 94.50% ($\downarrow$2.00%) 281.65 ($\uparrow$309.29) 0.38 ($\uparrow$0.01)
Table 4: Automatic evaluation results of counterfactuals generated by FitCF using Qwen2.5-72B, without label flip verification.
Table 4 divulges trends similar to those observed in Table 26.2.1). Omitting label flip verification leads to decreases in SLFR across both datasets, highlighting the importance of this step. However, skipping label flip verification occasionally results in lower PPL for certain methods, suggesting improved fluency in some cases.
Meanwhile, the decrease in SLFR is more pronounced for AG News, particularly with the Gradient method, which shows the largest SLFR drop alongside increases in PPL. Conversely, Integrated Gradients and LIME present minimal impact on SLFR, indicating a relative reliance on label flip verification to maintain consistent performance.
6.3 Discussion
Important words identified through feature attribution scores for BERT are more effective and less prone to hallucination for counterfactual generation compared to those self-generated by LLMs. Through ablation studies on the three core components of FitCF, we conclude that the number of demonstrations generated by ZeroCF has the most significant impact on the performance of FitCF. While specifying important words and applying label flip verification also contribute to FitCF’s effectiveness, their influence is less marked compared to the number of demonstrations. While SLFR decreases across three tables, the edit distance gets improved overall, except for SST, where no important words are specified. This indicates that without a certain component, the counterfactuals generated by FitCF are generally less edited, resulting in less successful label flips. Moreover, FitCF with Gradient proves to be the least robust, showing substantial drops in SLFR, when any of the three components is removed. In contrast, FitCF with LIME and SHAP demonstrate greater robustness and consistently produce high-quality counterfactuals.
Model Dataset AG News SST2
Method comp. suff. $\tau$ (loo) comp. suff. $\tau$ (loo)
Llama3 Gradient 0.20 0.13 0.06 0.21 0.25 -0.03
IG 0.38 0.03 0.07 -0.52 0.05 0.22
LIME 0.61 -0.02 0.16 0.68 0.02 0.29
SHAP 0.62 -0.02 0.16 0.60 0.03 0.25
Qwen-32B Gradient 0.12 0.12 0.07 0.20 0.23 -0.03
IG 0.32 0.03 0.05 0.50 0.04 0.21
LIME 0.53 -0.01 0.12 0.67 0.01 0.29
SHAP 0.53 -0.01 0.08 0.59 0.02 0.25
Qwen-72B Gradient 0.12 0.12 0.07 0.20 0.23 -0.03
IG 0.32 0.03 0.05 0.50 0.04 0.21
LIME 0.53 -0.01 0.12 0.67 0.01 0.29
SHAP 0.53 -0.01 0.07 0.59 0.02 0.25
Table 5: Faithfulness evaluation results based on Comprehensiveness (comp.), Sufficiency (suff.) and Kendall’s $\tau$ correlation with Leave-One-Out token removal ($\tau$ (loo)) for counterfactuals generated by FitCF using Llama3-8B, Qwen2.5-32B, and Qwen2.5-72B on AG News and SST2 datasets.
Figure 3: A Kendall’s tau ($\tau$) that quantifies the degree of correspondence between the ranking of generated counterfactuals’ quality and the ranking of feature attribution evaluation results is reported.
6.4 Correlation Analysis
From Table 5, we discover that LIME and SHAP consistently outperform Gradient and Integrated Gradients in terms of comprehensiveness and $\tau$ (loo) across all models and datasets, which aligns with our findings in §6.3. In addition, the comprehensiveness and sufficiency scores exhibit less variation across three models for AG News, though they are generally lower than those for SST2. In contrast, $\tau$ (loo) scores for SST2 are slightly higher compared to AG News. Furthermore, for AG News, a strong correlation ($\tau=1$) is observed in Figure 3 between the quality of generated counterfactuals and sufficiency, while for SST2, both comprehensive and $\tau$ (loo) demonstrate notable correlations with counterfactual quality. We conclude that the faithfulness of feature attribution scores is generally strongly correlated with the quality of counterfactuals generated with the auxiliary assistance of extracted important words using FitCF.
7 Conclusion
We first introduced ZeroCF, an approach that leverages important words derived from feature attribution methods for counterfactual example generation in a zero-shot setting. Building on this, we proposed FitCF, a framework that automatically constructs high-quality demonstrations using ZeroCF, eliminating the need for human-annotated ground truth for counterfactual generation. FitCF validates counterfactuals via label flip verification for their suitability as demonstrations in a few-shot setting. Empirically, FitCF outperforms three state-of-the-art baselines Polyjuice, BAE and FIZLE, and our own ZeroCF. Through ablation studies, we identified the three core components of FitCF - number of demonstrations, important words, and label flip verification - as critical to enhancing counterfactual quality. Moreover, we evaluated the faithfulness of feature attribution scores and found that LIME and Integrated Gradients are the most effective feature importance methods for FitCF, consistently producing the most faithful feature attribution scores. Finally, our analysis revealed a strong correlation between the faithfulness of feature attribution scores and the quality of the generated counterfactuals.
Future work includes investigating the correlation between additional dimensions of feature attribution scores, such as plausibility, coherence and insightfulness, and the quality of counterfactuals through user studies Domnich et al. (2025). We also plan to explore the potential of language models with architectures beyond encoder-only models as a foundation for feature attributions to be used in ZeroCF and FitCF. Furthermore, we would like to explore counterfactual example generation beyond typical text classification tasks to include, for instance, (long-form) question answering and open-ended mathematical problem, which are particularly challenging even for LLMs Dehghanighobadi et al. (2025).
Limitations
We conducted experiments exclusively using datasets in English. In other languages, the current approach may not offer the same advantages.
In ZeroCF and FitCF, feature attribution scores are determined by an explanation method for the predictions of a BERT model fine-tuned on the target dataset and the same BERT model is used to verify label flips. The potential contribution of other language models to performing both tasks in ZeroCF and FitCF, however, remains unexplored.
For feature attribution score evaluation, we focus only on faithfulness, although ferret also supports the evaluation of plausibility. However, plausibility evaluation requires human annotations, so we have left it out and consider it as future work.
Acknowledgment
We are indebted to the anonymous reviewers of ACL 2025 for their helpful and rigorous feedback. This work has been supported by the German Federal Ministry of Research, Technology and Space as part of the projects VERANDA (16KIS2047) and BIFOLD 24B.
Appendix A Prompt Instruction
A.1 Prompt for ZeroCF
You are an excellent assistant for text
editing. You are given an input from the
{dataset} dataset, classified into one of
{len(labels)} categories:
{’, ’.join(labels)}. The input belongs to
the ’{prediction}’ category.
{important_words} might be important words
leading to the ’{prediction}’ category.
Your task is to make minimal changes on the
below provided input to alter the
prediction category by carefully
considering provided important words.
Please output only the edited input.
Input: {input_text}
A.2 Prompt for FitCF
You are an excellent assistant for text
editing. You are given an input from the
{dataset} dataset, classified into one of
{len(labels)} categories:
{’, ’.join(labels)}. The input belongs to
the ’{prediction}’ category.
{important_words} might be important words
leading to the ’{prediction}’ category.
Your task is to make minimal changes on the
input provided below to alter the
prediction category to ’{counterpart}’ by
carefully considering provided important
words and examples. Please output the
edited input only!
Below are some examples consisting of
original and edited input.
[original input] {original_input_1}
[edit input] {edit_input_1}
[original input] {input_text}
[edit input]
Appendix B Detailed Information of Deployed BERT
Dataset Model Accuracy Link
AG News textattack/bert-base-uncased-ag-news 93.03% https://huggingface.co/textattack/bert-base-uncased-ag-news
SST2 gchhablani/bert-base-cased-finetuned-sst2 92.32% https://huggingface.co/gchhablani/bert-base-cased-finetuned-sst2
Table 6: BERT models used for AG News and SST2 datasets, with accuracy validated on their respective testsets.
Table 6 displays BERT models used for AG News and SST2 datasets with their validation accuracies. As both BERT models demonstrate strong performance in accuracy, we can use them as classifiers (§3.1) and label flip verifiers (§3.2).
Appendix C Visualization of Clustering
Figure 4 visualizes the clustering of sentence embeddings from AG News, and SST2 datasets, with their dimensions reduced to two using PCA. The illustrations suggest that generic patterns already exist, with instances from various clusters contributing to these patterns.
(a) AG News
(b) SST2
Figure 4: Visualization of clustering in AG News and SST2, where stars denote cluster centroids.
Appendix D Demonstration Selection by FitCF
Text Counterfactual
Bovina ends two-year wait. Seventh-seeded Russian Elena Bovina won her first title in two years by beating France’s Nathalie Dechy 6-2 2-6 7-5 in the final of the Pilot Pen tournament. Bovina ends two-year wait. Seventh-seeded Russian Elena Bovina won her first title in two years by beating France’s Nathalie Dechy 6-2 2-6 7-5 in the final of the International Event.
Wall St.’s Nest Egg - the Housing Sector NEW YORK (Reuters) - If there were any doubts that we’re still living in the era of the stay-at-home economy, the rows of empty seats at the Athens Olympics should help erase them. The Olympics - the Housing Sector NEW YORK (Reuters) - If there were any doubts that we’re still living in the era of the stay-at-home economy, the rows of empty seats at the Athens Olympics should help erase them.
French Take Gold, Bronze in Single Kayak ATHENS, Greece - Winning on whitewater runs in the family for Frenchman Benoit Peschier, though an Olympic gold is something new. Peschier paddled his one-man kayak aggressively but penalty free in both his semifinal and final runs on the manmade Olympic French Take Gold, Bronze in Single Kayaking Competition ATHENS, Greece - Winning on whitewater runs in the family for Frenchman Benoit Peschier, though an Olympic gold is something new. Peschier paddled his one-man kayak aggressively but without penalty in both his semifinal and final runs on the man-made Olympic course.
Japanese Utility Plans IPO in October (AP) AP - Electric Power Development Co., a former state-run utility, said Friday it is planning an initial public offering on the Tokyo Stock Exchange in October, a deal that could be the country’s biggest new stock listing in six years. Electric Power Development Co., a former state-run utility, is planning an initial public offering on the Tokyo Stock Exchange in October, a deal that could be the country’s biggest new stock listing in six years.
Afghan women make brief Olympic debut Afghan women made a short-lived debut in the Olympic Games on Wednesday as 18-year-old judo wildcard Friba Razayee was defeated after 45 seconds of her first match in the under-70kg middleweight. Afghan women make brief debut in international relations as 18-year-old Friba Razayee was defeated after 45 seconds of her first match in the under-70kg middleweight.
Table 7: The most similar demonstrations selected from each cluster for the question “Rivals Try to Turn Tables on Charles Schwab By MICHAEL LIEDTKE SAN FRANCISCO (AP) – With its low prices and iconoclastic attitude, discount stock broker Charles Schwab Corp. (SCH) represented an annoying stone in Wall Street’s wing-tipped shoes for decades…” from AG News. Corresponding counterfactuals are generated by Qwen2.5-72B using ZeroCF. Differences are marked in bold and edits are highlighted in red.
Table 7 shows the most similar demonstrations selected from each cluster, as shown in Figure 4 for the question “Rivals Try to Turn Tables on Charles Schwab By MICHAEL LIEDTKE SAN FRANCISCO (AP) – With its low prices and iconoclastic attitude, discount stock broker Charles Schwab Corp. (SCH) represented an annoying stone in Wall Street’s wing-tipped shoes for decades…” from AG News.
The decrease in SLFR performance while using a strong LLM can be attributed to the advanced contextual understanding of such models, e.g., Qwen2.5-72B. These models are more adept at discerning the underlying context of inputs and therefore less likely to incorrectly flip labels. For instance, as shown in Table 7, the second example remains clearly related to business, as the main topic—Housing Sector—is still evident, even though “Wall St.’s Nest Egg” is replaced with “The Olympic”.
Appendix E Dataset
E.1 Label Distribution
Figure 5 shows the label distributions of AG News and SST2 validation sets.
(a) AG News
(b) SST2
Figure 5: Label distribution of AG News and SST2.
E.2 Dataset Example
Figure 6 demonstrates example instances and gold labels from AG News and SST2 datasets.
Figure 6: Example instances from AG News and SST2.
Appendix F Experiment
F.1 Models
Table 8 demonstrates LLMs that are used for ZeroCF and FitCF. To reduce memory consumption, we use a GPTQ-quantized version Frantar et al. (2023). All LLMs are directly downloaded from Huggingface and run on a single NVIDIA RTXA6000, A100 or H100 GPU.
Table 8: Three open sourced LLMs used in ZeroCF and FitCF.
F.2 Inference Time
AG News SST2
ZeroCF FitCF ZeroCF FitCF
Llama3-8B 8h 13h 2h 5h
Qwen2.5-32B 9h 17h 7h 12h
Qwen2.5-72B 38h 47h 8h 16h
Table 9: Inference time for ZeroCF and FitCF using Llama3-8B, Qwen2.5-32B and Qwen2.5-32B on AG News and SST2.
Table 9 shows inference time for ZeroCF and FitCF using Llama3-8B, Qwen2.5-32B and Qwen2.5-32B across AG News and SST2 datasets.
Appendix G Calculation of Label Flip Rate
We use the same LLM to serve as both the flip label verifier and the counterfactual generator (§5.1). To validate deployed LLMs’ classification performance, we evaluate them on the AG News and SST2 datasets. Subsequently, we detail the prompt instructions used for flip label verification.
G.1 Classification Performance of LLMs
Dataset Model Accuracy Error Rate
AG News Llama3-8B 72.39% 0.70%
Qwen2.5-32B 80.73% 0.28%
Qwen2.5-72B 79.12% 0.47%
SST2 Llama3-8B 89.75% 0.00%
Qwen2.5-32B 94.61% 0.00%
Qwen2.5-72B 94.27% 0.11%
Table 10: Accuracy score and error rate on AG News and SST2 datasets across three runs on the validation set using Llama3-8B, Qwen2.5-32B, and Qwen2.5-72B in a zero-shot setting. The error rate is calculated by counting the number of instances where the predicted label falls outside the pre-defined label set.
Table 10 displays the accuracy score and error rate on AG News and SST2 datasets using Llama3-8B, Qwen2.5-32B, and Qwen2.5-72B. Our findings indicate that Qwen2.5-32B demonstrates the best classification performance with the lowest error rate, whereas Llama3-8B has the poorest classification performance. Notably, Qwen2.5-72B is the only LLM that generates predictions outside the predefined labels on SST2. Moreover, as selected LLMs and BERT perform similarly on the two datasets (Table 6), we can assume that the tendency of SLFR will be consistent, and our conclusion should remain unchanged.
G.2 Prompt Instruction
You are an excellent assistant for text
classification. You are provided with an
original and an edited instance from the
{dataset_name} dataset. Each instance
belongs to one of {len(labels)} categories:
{’, ’.join(labels)}. Determine if the
predicted classifications of the original
and edited instances are different.
[original instance] ’{instance}’
[edited instance] ’{counterfactual}’
Respond with ’yes’ if they are different.
Response with ’no’ if they are the same.
Answer ’yes’ or ’no’ only!